# Absolute continuity of non-Gaussian and Gaussian multiplicative chaos measures

### v1

@@ 1.3. Outline of the proof and structure of the paper
The discretization step proceeds by considering a nested sequence of partitions of $[0,1]$ such that the $n^{\mathrm{th}}$ partition $\mathcal{P}_{n}$ contains approximately $2^{n}$ intervals (we’ll need slightly more, see (2.1)). We then define the random series $\widetilde{S}_{n,a}(t)$, defined in Eq. 2.6, which is a version of the random Fourier series $S_{2^{n-1},a}(t)$ that is (i) composed of dyadically-growing chunks of $S_{2^{n-1},a}(t)$, so that the $\ell^{\mathrm{th}}$ increment of $\widetilde{S}_{n,a}(t)$, $\ell\in\llbracket 1,n\rrbracket$, is composed of a sum of $\approx 2^{\ell}$ terms of $S_{2^{n-1},a}$ and therefore has approximately constant variance $\log 2$ for all $\ell$; and (ii) is constant on each interval in $\mathcal{P}_{n}$. In light of (ii) and the fact that the $\mathcal{P}_{n}$ are nested, the field $(\widetilde{S}_{n,a}(t))_{t\in[0,1]}$ may be defined on a tree that is approximately binary; as such, the field resembles an inhomogeneous branching random walk (the major difference being that particles do not move independently after splitting). Indeed, this is the perspective we take, and we write $(\widetilde{S}_{n,a}(v))_{v\in\mathcal{N}_{n}}\equiv(\widetilde{S}_{n,a}(t))_{t\in[0,1]}$, where $\mathcal{N}_{n}$ denotes the set of particles in the $n^{\mathrm{th}}$ generation of the tree. Using the regularity of the eigenbasis $\{\cos(2\pi kt),\sin(2\pi kt)\}_{k\in\mathbb{N}}$, we show in Proposition 2.3 via a chaining argument that the chaos measure $\widetilde{\mu}_{\gamma,\widetilde{a}}$ associated to $(\widetilde{S}_{n,a}(v))_{v\in\mathcal{N}_{n},n\in\mathbb{N}}$ is mutually absolutely continuous with respect to $\mu_{\gamma,a}$, almost surely.
@@
The coupling step proceeds as follows. Note by (i) above and the central limit theorem that the $n^{\mathrm{th}}$ increment $\widetilde{a}_{n}(v)$ of $\widetilde{S}_{n,a}(v)$ becomes more and more Gaussian as $n$ increases. However, the increments $\{\widetilde{a}_{n}(v)\}_{v\in\mathcal{N}_{n}}$ are not independent from one another, and therefore we cannot couple them to Gaussian variables one-by-one. To circumvent this, we use the coupling result of Yurinskii (Lemma 2.4 here, recorded from [BELLONI20194] and originating from [yurinskii1978error]) to directly couple a well-chosen subset of $\{\tilde{a}_{n}(v)\}_{v\in\mathcal{N}_{n}}$ as a vector to a Gaussian vector with the same covariance structure. While the coupling error provided by Yurinskii is insufficient for coupling the entire vector of increments at level $n$ to be to close (in $L^{\infty}$-norm) to a Gaussian vector, the critical insight is that we only need to closely couple the subset of increments which are descendants of “$\gamma$-thick points” of the field. Thus, we do not need to couple a vector of length $|\mathcal{P}_{n}|\approx 2^{n}$ but instead make do with a vector of length $\approx 2^{q_{\gamma}n}$, for some $q_{\gamma}$ that tends to 00 as $\gamma\to\sqrt{2}$. This is accomplished in Lemma 2.6. Note the requirement $\gamma>1$ is introduced to ensure that the amount of random variables we need to closely couple is small enough and can be read explicitly from the right-hand side of (2.20).
@@
Ultimately, this leads to a coupled Gaussian field $\widetilde{S}_{g}:=(\widetilde{S}_{n,g}(v))_{v\in\mathcal{N}_{n},n\in\mathbb{N}}$ constructed on the same tree as $(\widetilde{S}_{n,a}(v))_{v\in\mathcal{N}_{n},n\in\mathbb{N}}$. Proposition 2.7 states that their associated chaos measures are mutually absolutely continuous. Up to and including this result, we have shown that $\mu_{\gamma,a}$ is absolutely continuous with respect to the GMC associated to the discrete, hierarchical Gaussian model $\widetilde{S}$. The proof of 1.1, given in Section 2.3, follows because the construction of the hierarchical model is in a sense “invertible”: an associated Fourier series with i.i.d. $\mathcal{N}(0,1)$ coefficients can be constructed from $\widetilde{S}_{g}$.
@@
The proof of Proposition 2.7 occupies the bulk of the paper. We construct the Radon-Nikodym derivative between these two chaos measures by studying integrability properties of the obvious candidate, defined in (3.1), with respect to each chaos measure. Our proof proceeds via a thick-point analysis and moment estimates using barriers. We refer to the start of Section 3 for a more technical outline.
@@
In Section 2, we carry out the discretization and coupling steps. In Section 3, we prove Proposition 2.7. In Section 4, we discuss open questions raised by our work, as well as the extension to general bounded domains (see Theorem 4.1).
@@ 1.4. Some notation
Below, we use $c,C>0$ to denote constants depending only on the distribution of $a_{1}^{(1)}$ that may change from line-to-line. We also say $f(n)=O_{n}(g(n))$ if $|f(n)|\leq Cg(n)$ for all $n\geq 1$, and $f(n)=o_{n}(g(n))$ if $f/g\to 0$ as $n\to\infty$. The set of positive integers is denoted by $\mathbb{N}$.
2. Discretization, Gaussian coupling, and proof of 1.1
In Section 2.1, we define a discretization of the field (1.1) that lives on a hierarchical lattice. We then state and prove Proposition 2.3, which states absolute continuity of $\mu_{\gamma,a}$ with the chaos measure associated to this discrete hierarchical model. In Section 2.2, we construct a coupling between the hierarchical model and a Gaussian hierarchical model using the Yurinskii coupling [yurinskii1978error] in a version for the supremum norm proved in [BELLONI20194]. We then state Proposition 2.7, which yields absolute continuity between the associated chaos measures. In Section 2.3, we prove 1.1.
@@ 2.1. The discrete hierarchical model
We begin by defining a sequence of partitions $\{\mathcal{P}_{n}\}_{n\geq 1}$ of $[0,1]$ in the following manner. Let $f:\mathbb{N}\to\mathbb{R},n\mapsto n^{4}$ and $\mathcal{S}_{1}:=\{0,1\}$. Given $\mathcal{S}_{n}$ we define
@@
At this point, it is useful to record
@@
For each $n\geq 1$ and $j\in\llbracket 1,T_{n}\rrbracket$, we choose a representative element $t_{n}(j)$ of $I_{n}(j)$ in such a way that every $2^{n-1}\times 2^{n-1}$ square submatrix of
@@
is invertible. This is possible because the zero-set of any non-zero analytic function from $\mathbb{R}^{N}\to\mathbb{R}$ has measure 00, and the exact choice of $t_{n}(j)$ is of no importance. (This invertibility condition allows us to construct a Fourier series with i.i.d. random coefficients from a discrete hierarchical model, which leads to the construction of the $g_{k}^{(i)}$ appearing in 1.1, see (2.23).)
@@
The following notation will make this bijection explicit, translating between the tree $\mathfrak{T}$ and the system of representative elements. We let $\mathcal{N}_{n}$ denote the collection of vertices in the $n^{\mathrm{th}}$ level of $\mathfrak{T}$, so that $|\mathcal{N}_{n}|=T_{n}$ and $V(\mathfrak{T})=\cup_{n\geq 1}\mathcal{N}_{n}$. For $v\in\mathcal{N}_{n}$ corresponding to $t_{n}(j)$, define
@@
(i.e., $t_{m}(v)$ is the representative of the generation-$m$ parent interval corresponding to $v$). For each $n\geq 1$, let
@@
$C_{k,n}^{(1)}(v)$ $:=\frac{1}{\sqrt{k}}\cos(2\pi kt_{n}(v))\quad\text{and}\quad C_{k,n}^{(2)}(v):=\frac{1}{\sqrt{k}}\sin(2\pi kt_{n}(v))\,,$ (2.3)
$\widetilde{a}_{n}^{(i)}(v)$ $:=(C_{k,n}^{(i)}(v))_{k\in\{2^{n-1},\dots,2^{n}-1\}}\cdot(a^{(i)}_{k})_{k\in\{2^{n-1},\dots,2^{n}-1\}}\,,\text{ for }i\in\{1,2\}\,,\text{ and}$
$\widetilde{a}_{n}$ $:=\widetilde{a}_{n}^{(1)}+\widetilde{a}_{n}^{(2)}.$
Observe that our choice of the $t_{n}(v)$ ensures that any $2^{n-1}\times 2^{n-1}$-submatrix of $(C_{k,n}^{(i)}(v))_{k,v}$ is invertible (see below Eq. 2.2). Observe further that $\widetilde{a}_{n}$ is a vector in $\mathbb{R}^{T_{n}}$, recalling $T_{n}=\mathcal{N}_{n}$. Denote its coordinates by $\widetilde{a}_{n}=:(\widetilde{a}_{n}(v))_{v\in\mathcal{N}_{n}}$, and note (2.3) gives
@@
Though we prefer to think of our model as living on a tree, it will occasionally be more convenient for calculations to define
$C_{k,n}^{(i)}(t):=C_{k,n}^{(i)}(v_{n}(t))\qquad\text{and}\qquad\widetilde{a}_{n}(t):=\widetilde{a}_{n}(v_{n}(t))\,.$
@@
$\widetilde{S}_{k,a}(v):=\sum_{m=1}^{k}\widetilde{a}_{m}(v)\qquad\text{and}\qquad\widetilde{S}_{k,a}(t):=\sum_{m=1}^{k}\widetilde{a}_{m}(v_{m}(t))\,,\text{ for }1\leq k\leq n\,.$ (2.6)
@@
Finally, we are ready to define the associated chaos measure. For $n\in\mathbb{N}$, $\gamma\in\mathbb{R}$, $(a^{(i)}_{k})_{k\in\mathbb{N},i\in\{1,2\}}$ i.i.d. satisfying (1.2), and $A\subset[0,1]$ Borel, we set
@@
where $\widetilde{Z}_{n,\gamma,\widetilde{a}}(t):=\mathbb{E}[e^{\gamma\widetilde{S}_{n,a}(v_{n}(t))}]$. In Remark 2.1, we show that $\widetilde{S}_{n,\widetilde{a}}(t)$ is a log-correlated field so that [Junnila, Theorem 2] implies $\widetilde{\mu}_{n,\gamma,\widetilde{a}}$ converges to a random measure $\widetilde{\mu}_{\gamma,\widetilde{a}}$, which we call the (discrete) hierarchical chaos measure corresponding to $(a_{k}^{(i)})_{k\in\mathbb{N},i\in\{1,2\}}$.
@@
The fields $(S_{n,a}(t))_{t\in[0,1]}$ and $(\widetilde{S}_{n,a}(t))_{t\in[0,1]}$, defined in Eqs. 1.1 and 2.6 respectively, are log-correlated fields. In particular, their increments fulfill the conditions in [Junnila, Definition 4]. Indeed, that $(S_{n,a}(t))_{t\in[0,1]}$ satisfies these conditions is shown in [Junnila, Section 3.1], and using the identity Eq. 2.7, it follows immediately that $(\widetilde{S}_{n,a}(t))_{t\in[0,1]}$ does as well. Therefore, we may apply the estimates developed in [Junnila] for such fields to $S_{n,a}$ and $\widetilde{S}_{n,a}$. We repeatedly use the following result on the Laplace transform of $\widetilde{S}_{n,\widetilde{a}}(t)$: [^1] As written, [Junnila, Lemma 7] collects the $\widetilde{\zeta}(\gamma,t)+O_{n}(2^{-n/2})$ term into a term simply denoted $\epsilon$. Our formulas for this $\epsilon$ can be deduced from the final two displays of the proof of [Junnila, Lemma 7], taking their $\xi_{i}$ to be our $\gamma_{i}$, their $\varphi(\gamma_{1},\gamma_{2})$ to be $\mathbb{E}[e^{\gamma_{1}\widetilde{S}_{n,\widetilde{a}}(t)+\gamma_{2}\widetilde{S}_{n,\widetilde{a}}(s)}]$, and absorbing the terms in the final display of the proof of [Junnila, Lemma 7] depending only on $k_{0}$ into our $\widetilde{\zeta}(\gamma_{1},\gamma_{2},t,s)$. The summation over $d_{k}O(|\xi|^{3}+|\xi|^{6})$ term in the final display of the proof of [Junnila, Lemma 7] converges absolutely because $\sum_{k}d_{k}<\infty$. Their proof in fact gives $d_{k}\leq O_{k}(k^{-3/2})$ due to the $k^{-1/2}$ term present in our $X_{k}$ and $\widetilde{a}_{k}$ (the $d_{k}$ are dominated by a $(3+\varepsilon)$ moment bound), so that $\sum_{k\geq 2^{n}-1}d_{k}\leq O_{n}(2^{-n/2})$.
Lemma 2.2 ([Junnila, Lemma 7]).
@@
$\mathbb{E}[e^{\gamma_{1}\widetilde{S}_{n,\widetilde{a}}(t)+\gamma_{2}\widetilde{S}_{n,\widetilde{a}}(s)}]=e^{\frac{\gamma_{1}^{2}}{2}\operatorname{Var}[\widetilde{S}_{n,\widetilde{a}}(t)]+\frac{\gamma_{2}^{2}}{2}\operatorname{Var}[\widetilde{S}_{n,\widetilde{a}}(s)]+\gamma_{1}\gamma_{2}\operatorname{Cov}[\widetilde{S}_{n,\widetilde{a}}(t),\widetilde{S}_{n,\widetilde{a}}(s)]+\widetilde{\zeta}(\gamma_{1},\gamma_{2},t,s)+O_{n}(2^{-n/2})},$
where $|\zeta(\gamma_{1},\gamma_{2},t,s)|\leq C_{R}$ is bounded by a constant depending only on $R$ and the distribution of $a$.

### v2

@@ 1.3. Outline of the proof and structure of the paper
The discretization step proceeds by considering a nested sequence of partitions of $[0,1]$ such that the $n^{\mathrm{th}}$ partition $\mathcal{P}_{n}$ contains approximately $2^{n}$ intervals (we will need slightly more, see Eq.˜2.1). We then define the random series $\widetilde{S}_{n,a}(t)$ in Eq.˜2.6 which is a modification of the random Fourier series $S_{2^{n-1},a}(t)$. The main characteristics of $\widetilde{S}_{n,a}(t)$ are that it (i) is composed of dyadically-growing chunks of $S_{2^{n-1},a}(t)$, so that the $\ell^{\mathrm{th}}$ increment of $\widetilde{S}_{n,a}(t)$, $\ell\in\llbracket 1,n\rrbracket$, is composed of a sum of $\approx 2^{\ell}$ terms of $S_{2^{n-1},a}$ and therefore has approximately constant variance $\log 2$ for all $\ell$; and (ii) is constant on each interval in $\mathcal{P}_{n}$. In light of (ii) and the fact that the $\mathcal{P}_{n}$ are nested, the field $(\widetilde{S}_{n,a}(t))_{t\in[0,1]}$ may be defined on an approximately binary tree; as such, the field resembles an inhomogeneous branching random walk (the major difference being that particles do not move independently after splitting). Indeed, this is the perspective we take, and we write $(\widetilde{S}_{n,a}(v))_{v\in\mathcal{N}_{n}}\equiv(\widetilde{S}_{n,a}(t))_{t\in[0,1]}$, where $\mathcal{N}_{n}$ denotes the set of particles in the $n^{\mathrm{th}}$ generation of the tree. Using the regularity of the eigenbasis $\{\cos(2\pi kt),\sin(2\pi kt)\}_{k\in\mathbb{N}}$, we show in Proposition˜2.3 via a chaining argument that the chaos measure $\widetilde{\mu}_{\gamma,\widetilde{a}}$obtained from $(\widetilde{S}_{n,a}(v))_{v\in\mathcal{N}_{n},n\in\mathbb{N}}$ by replacing $S$ by $\widetilde{S}$ in (1.3) is mutually absolutely continuous with respect to $\mu_{\gamma,a}$, almost surely.
@@
The coupling step proceeds as follows. Note that by (i) above and the central limit theorem the $n^{\mathrm{th}}$ increment $\widetilde{a}_{n}(v)$ of $\widetilde{S}_{n,a}(v)$ becomes more and more Gaussian as $n$ increases. However, the increments $\{\widetilde{a}_{n}(v)\}_{v\in\mathcal{N}_{n}}$ are not independent from one another, and therefore we cannot couple them to Gaussian variables one-by-one. To circumvent this, we use the coupling result of Yurinskii (Lemma˜2.4 here, recorded from [5] and originating from [45]) to directly couple a well-chosen subset of $\{\tilde{a}_{n}(v)\}_{v\in\mathcal{N}_{n}}$ as a vector to a Gaussian vector with the same covariance structure. A naive application of the Yurinskii coupling—i.e., trying to couple the entire vector of increments at level $n$ such that it is close in $L^{\infty}$-norm to a Gaussian vector—will fail because the dimension of this vector is too high. The critical insight to circumvent this issue is that we will only need to closely couple the subset of increments which are descendants of $\gamma$-thick points” of the field. Thus, we do not need to couple a vector of length $|\mathcal{P}_{n}|\approx 2^{n}$ but instead make do with a vector of length $\approx 2^{q_{\gamma}n}$, for some $q_{\gamma}$ that tends to 0 as $\gamma\to\sqrt{2}$. This is accomplished in Lemma˜2.6. Note the requirement $\gamma>1$ is introduced to ensure that the amount of random variables we need to closely couple is small enough and can be read explicitly from the right-hand side of Eq.˜2.20.
@@
Ultimately, this leads to a coupled Gaussian field $\widetilde{S}_{g}:=(\widetilde{S}_{n,g}(v))_{v\in\mathcal{N}_{n},n\in\mathbb{N}}$ constructed on the same tree as $(\widetilde{S}_{n,a}(v))_{v\in\mathcal{N}_{n},n\in\mathbb{N}}$. Proposition˜2.7 states that their associated chaos measures are mutually absolutely continuous. Up to and including this result, we have shown that $\mu_{\gamma,a}$ is absolutely continuous with respect to the GMC associated to the discrete, hierarchical Gaussian model $\widetilde{S}$. The proof of ˜1.1, given in Section˜2.3, follows because the construction of the hierarchical model is in a sense “invertible”: an associated Fourier series with i.i.d. $\mathcal{N}(0,1)$ coefficients can be constructed from $\widetilde{S}_{g}$.
@@
The proof of Proposition˜2.7 occupies the bulk of the paper. We construct the Radon-Nikodym derivative between these two chaos measures by studying integrability properties of the obvious candidate, defined in (3.1), with respect to each chaos measure. Our proof proceeds via a thick-point analysis and moment estimates using barriers. We refer to the start of Section˜3 for a more technical outline.
@@
In Section˜2, we carry out the discretization and coupling steps. In Section˜3, we prove Proposition˜2.7. In Section˜4, we discuss open questions raised by our work, as well as the extension to general bounded domains (see Theorem˜4.1).
@@ 1.4. Some notation
Below, we use $c,C>0$ to denote constants depending only on the distribution of $a_{1}^{(1)}$ that may change from line to line. We also say $f(n)=O_{n}(g(n))$ if $|f(n)|\leq Cg(n)$ for all $n\geq 1$, and $f(n)=o_{n}(g(n))$ if $f/g\to 0$ as $n\to\infty$. The set of positive integers is denoted by $\mathbb{N}$. For $m\leq n\in\mathbb{N}$ we set $\llbracket m,n\rrbracket:=\{m,\dots,n\}$.
2. Discretization, Gaussian coupling, and proof of ˜1.1
In Section˜2.1, we define a discretization of the field (1.1) on a hierarchical lattice. We then state and prove Proposition˜2.3, which concerns mutual absolute continuity of $\mu_{\gamma,a}$ with the chaos measure associated to this discrete hierarchical model. In Section˜2.2, we construct a coupling between the hierarchical model and a Gaussian hierarchical model using the Yurinskii coupling [45] in a version for the supremum norm proved in [5]. We then state Proposition˜2.7, which yields absolute continuity between the associated chaos measures. In Section˜2.3, we prove ˜1.1.
@@ 2.1. The discrete hierarchical model
We begin by defining a sequence of partitions $\{\mathcal{P}_{n}\}_{n\geq 1}$ of $[0,1)$ in the following manner. Let $f:\mathbb{N}\to\mathbb{R},n\mapsto n^{4}$ and $\mathcal{S}_{1}:=\{0,1\}$. Given $\mathcal{S}_{n}$ we define
@@
We emphasize that the points $\tau_{n}(j)$ are not equally spaced, and so the intervals $I_{n}(j)$ do not necessarily have the same length. At this point, it is useful to record
@@
For each $n\geq 1$ and $j\in\llbracket 1,T_{n}\rrbracket$, we choose a representative element $t_{n}(j)$ of $I_{n}(j)$ in such a way that every $2^{n-1}\times 2^{n-1}$ square submatrix of both
@@
is invertible. This is possible because the determinants of the $2^{n-1}\times 2^{n-1}$ square submatrices of $C_{n}^{(1)}$ and $C_{n}^{(2)}$ are analytic functions of $(t_{n}(j))_{j\leq T_{n}}$; as such they have a zero set of Lebesgue measure 0, which implies that we can choose the $t_{n}(j)$ such that all these determinants are non-zero. The exact choice of $t_{n}(j)$ is of no importance. This invertibility condition allows us to construct a Fourier series with i.i.d. random coefficients from a discrete hierarchical model, which leads to the construction of the $g_{k}^{(i)}$ appearing in ˜1.1, see (2.23).
@@
The following notation makes this bijection explicit, translating between the tree $\mathfrak{T}$ and the system of representative elements. We let $\mathcal{N}_{n}$ denote the collection of vertices in the $n^{\mathrm{th}}$ level of $\mathfrak{T}$, so that $|\mathcal{N}_{n}|=T_{n}$ and $V(\mathfrak{T})=\cup_{n\geq 1}\mathcal{N}_{n}$. For $v\in\mathcal{N}_{n}$ corresponding to $t_{n}(j)$, define
@@
i.e., $t_{m}(v)$ is the representative of the generation-$m$ parent interval corresponding to $v$. For each $n\geq 1$, let
@@
$C_{k,n}^{(1)}(v)$ $=\frac{1}{\sqrt{k}}\cos(2\pi kt_{n}(v))\quad\text{and}\quad C_{k,n}^{(2)}(v)=\frac{1}{\sqrt{k}}\sin(2\pi kt_{n}(v))\,$, (2.3)
$\widetilde{a}_{n}^{(i)}(v)$ $=(C_{k,n}^{(i)}(v))_{k\in\{2^{n-1},\dots,2^{n}-1\}}\cdot(a^{(i)}_{k})_{k\in\{2^{n-1},\dots,2^{n}-1\}}\,,\text{ for }i\in\{1,2\}\,,\text{ and}$
$\widetilde{a}_{n}$ $=\widetilde{a}_{n}^{(1)}+\widetilde{a}_{n}^{(2)}.$
Observe that our choice of the $t_{n}(v)$ ensures that any $2^{n-1}\times 2^{n-1}$-submatrix of $(C_{k,n}^{(i)}(v))_{k,v}$ is invertible (see below Eq.˜2.2). Observe further that $\widetilde{a}_{n}$ is a vector in $\mathbb{R}^{T_{n}}$, recalling $T_{n}=|\mathcal{N}_{n}|$. Denote its coordinates by $\widetilde{a}_{n}:=(\widetilde{a}_{n}(v))_{v\in\mathcal{N}_{n}}$, and note (2.3) gives
@@
Though we prefer to think of our model as being defined on a tree, it will occasionally be more convenient for calculations to define
$C_{k,n}^{(i)}(t):=C_{k,n}^{(i)}(v_{n}(t))\qquad\text{and}\qquad\widetilde{a}_{n}(t):=\widetilde{a}_{n}(v_{n}(t))\quad\text{for}\ t\in[0,1]\,.$
@@
$\widetilde{S}_{k,a}(v):=\sum_{m=1}^{k}\widetilde{a}_{m}(v)\quad\text{and}\quad\widetilde{S}_{k,a}(t):=\sum_{m=1}^{k}\widetilde{a}_{m}(v_{m}(t))\,,\quad\text{ for }1\leq k\leq n\,.$ (2.6)
@@
Finally, we are ready to define the associated chaos measure. For $n\in\mathbb{N}$, $\gamma\in\mathbb{R}$, $(a^{(i)}_{k})_{k\in\mathbb{N},i\in\{1,2\}}$ i.i.d. satisfying Eq.˜1.2, and $A\subset[0,1]$ Borel, we set
@@
where $\widetilde{Z}_{n,\gamma,\widetilde{a}}(t):=\mathbb{E}[e^{\gamma\widetilde{S}_{n,a}(v_{n}(t))}]$. In Remark 2.1, we show that $\widetilde{S}_{n,\widetilde{a}}(t)$ is a log-correlated field so that [25, Theorem 2] implies $\widetilde{\mu}_{n,\gamma,\widetilde{a}}$ converges almost surely in the space of Radon measures on $[0,1]$ equipped with the topology of weak convergence to a random measure $\widetilde{\mu}_{\gamma,\widetilde{a}}$, which we call the hierarchical chaos measure corresponding to $(a_{k}^{(i)})_{k\in\mathbb{N},i\in\{1,2\}}$.
@@
The fields $(S_{n,a}(t))_{t\in[0,1]}$ and $(\widetilde{S}_{n,a}(t))_{t\in[0,1]}$, defined in Eqs.˜1.1 and 2.6 respectively, are log-correlated fields. In particular, their increments $X_{k}(t)$ fulfill the conditions in [25, Definition 4]. Indeed, that $(S_{n,a}(t))_{t\in[0,1]}$ satisfies these conditions is shown in [25, Section 3.1], and using the identity Eq.˜2.7, it follows immediately that $(\widetilde{S}_{n,a}(t))_{t\in[0,1]}$ does as well. Therefore, we may apply the estimates developed in [25] for such fields to $S_{n,a}$ and $\widetilde{S}_{n,a}$.
We repeatedly use the following result on the Laplace transforms of $S_{n,a}(t)$, $\widetilde{S}_{n,\widetilde{a}}(t)$:
Lemma 2.2.
@@
[…]
and
[…]
Proof.
The result essentially comes from examining the proof of [25, Lemma 7]. We shall focus on proving the second equality above, as the first equality follows similarly.
As written, [25, Lemma 7] [^1] For ease of reading, let us translate the notation appearing in [25, Lemma 7] and its proof to our notation. Take their $X_{k}$ to be our $X_{k}$, and take $j=n$, $t_{n}=2^{n}$, $x=\pi_{m}(t)$, $y=\pi_{m}(s)$, $\xi_{1}=\gamma_{1}$, and $\xi_{2}=\gamma_{2}$. collects the desired $\widetilde{\zeta}(\gamma_{1},\gamma_{2},t,s)+O_{n}(2^{-n/2})$ term into a term simply denoted by $\epsilon$, which is shown to be bounded by a constant $C_{R}$ uniformly over all $n,t,s,\gamma_{1}$, and $\gamma_{2}$. Therefore, it remains to extract this quantitative decomposition From $\epsilon$. The key estimate for this is the following: for any $p,q>0$, there exists a constant $\mathcal{E}(p+q)>0$ such that for all $t,s,\in[0,1]$ and $k\in\mathbb{N}$,
$\mathbb{E}\Big[|X_{k}(t)|^{p}|X_{k}(s)|^{q}\Big]\leq\mathcal{E}(p+q)k^{-\frac{p+q}{2}}\,.$
This is immediate from (1.2) and (2.5).
Now, as in the first line of the proof of [25, Lemma 7], define
$\varphi_{k}(\gamma_{1},\gamma_{2}):=\varphi_{k}(\gamma_{1},\gamma_{2},t,s)=\mathbb{E}[e^{\gamma_{1}X_{k}(\pi_{m}(t))+\gamma_{2}X_{k}(\pi_{m}(s))}]$
where $m=\lfloor\log_{2}k\rfloor+1$ and we recall $\pi_{m}$ from (2.1) (we will not use the precise definition of $\pi_{m}$ as all bounds will be given uniformly over $t$ and $s$). We then follow the proof of [25, Lemma 7] until the second-to-last display, where we instead define
$\varphi(\gamma_{1},\gamma_{2}):=\varphi(\gamma_{1},\gamma_{2},t,s)=\mathbb{E}[e^{\gamma_{1}\widetilde{S}_{n,\widetilde{a}}(t)+\gamma_{2}\widetilde{S}_{n,\widetilde{a}}(s)}]=\prod_{m=1}^{n}\prod_{k=2^{m-1}}^{2^{m}-1}\varphi_{k}(\gamma_{1},\gamma_{2},t,s)\,.$
Recall $k_{0}$ from the first line of [25, Page 6176], which was shown to depend only on $R$. Define $m_{0}:=\inf\{m\in\mathbb{N}:2^{m}-1\geq k_{0}\}$, and write
$\varphi(\gamma_{1},\gamma_{2})=\left(\prod_{m=1}^{m_{0}}\prod_{k=2^{m-1}}^{2^{m}-1}\varphi_{k}(\gamma_{1},\gamma_{2},t,s)\right)\left(\prod_{m=m_{0}+1}^{n}\prod_{k=2^{m-1}}^{2^{m}-1}\varphi_{k}(\gamma_{1},\gamma_{2},t,s)\right)$
(this takes the place of the second-to-last display of the proof of [25, Lemma 7]). Denote the logarithm of the first product by
$\widetilde{\zeta}_{1}(\gamma_{1},\gamma_{2},t,s):=\log\left(\prod_{m=1}^{m_{0}}\prod_{k=2^{m-1}}^{2^{m}-1}\varphi_{k}(\gamma_{1},\gamma_{2},t,s)\right)\,$,
which for all $\gamma_{1},\gamma_{2},t$ and $s$ is bounded by a constant depending only on $R$ (due to the same being true for $m_{0}$ and (1.2)) and will eventually be absorbed into $\widetilde{\zeta}(\gamma_{1},\gamma_{2},t,s)$. Similar to the final display of [25, Lemma 7], the logarithm of the second product is
$\frac{\gamma_{1}^{2}}{2}\operatorname{Var}[\widetilde{S}_{n,\widetilde{a}}(t)]+\frac{\gamma_{2}^{2}}{2}\operatorname{Var}[\widetilde{S}_{n,\widetilde{a}}(s)]+\gamma_{1}\gamma_{2}\operatorname{Cov}[\widetilde{S}_{n,\widetilde{a}}(t),\widetilde{S}_{n,\widetilde{a}}(s)]$
$+\widetilde{\zeta}_{2}(\gamma_{1},\gamma_{2},t,s)-\sum_{k=2^{n}}^{\infty}d_{k}O(\gamma^{3}+\gamma^{6})\,$,
where $\gamma:=\max(|\gamma_{1}|,|\gamma_{2}|)$, the $O$ term is independent of $t$ and $s$, and
$\widetilde{\zeta}_{2}(\gamma_{1},\gamma_{2},t,s)$
$:=$ $-\frac{\gamma_{1}^{2}}{2}\sum_{k=1}^{2^{m_{0}}-1}\mathbb{E}[X_{k}(\pi_{m}(t))^{2}]-\frac{\gamma_{2}^{2}}{2}\sum_{k=1}^{2^{m_{0}}-1}\mathbb{E}[X_{k}(\pi_{m}(s))^{2}]$
$-\gamma_{1}\gamma_{2}\sum_{k=1}^{2^{m_{0}}-1}\mathbb{E}[X_{k}(\pi_{m}(t))X_{k}(\pi_{m}(s))]+\sum_{k=2^{m_{0}}-1}^{\infty}d_{k}O(|\gamma|^{3}+|\gamma|^{6})\,.$
The equations leading up to [25, Eq.(14)] and our moment bound on $\mathbb{E}[|X_{k}(t)|^{p}|X_{k}(s)|^{q}]$ in the first display of the proof show that $|c_{k}|=O(k^{-3/2})$. Our moment bound and the definition of $d_{k}$ in [25] imply $|d_{k}|=(1+o(1))|c_{k}|=O(k^{-3/2})$. Therefore, we find $|\widetilde{\zeta}_{2}(\gamma_{1},\gamma_{2},t,s)|$ is bounded by a constant depending only on $R$, and
$\sum_{k=2^{n}}^{\infty}d_{k}O(\gamma^{3}+\gamma^{6})=O(2^{-n/2})\,.$
Taking $\widetilde{\zeta}=\widetilde{\zeta}_{1}+\widetilde{\zeta}_{2}$ concludes the proof. ∎